The sum of the three numbers 138. the second is 5 more than the smallest and the third is 10 more than the smallest.
Answer by stanbon(48558)

The sum of the three numbers 138. the second is 5 more than the smallest and the third is 10 more than the smallest.
----------------------------------------------
Let "x" be the smallest:
x + y + z = 138
y = x+5
z = x+10
---------------
Substitute to get:
x + x+5 + x+10 = 138
3x + 15 = 138
3x = 123
x = 123/3
y = (123/3)+5
z = (1/23/3) + 10
===========================
Cheers,
Stan H.